### v1

# Towards Multi-agent Policy-based Directed Hypergraph Learning for Traffic Signal Control

Abstract
Deep reinforcement learning (DRL) methods that incorporate graph neural networks (GNNs) have been extensively studied for intelligent traffic signal control, which aims to coordinate traffic signals effectively across multiple intersections. Despite this progress, the standard graph learning used in these methods still struggles to capture higher-order correlations in real-world traffic flow. In this paper, we propose a multi-agent proximal policy optimization framework DHG-PPO, which incorporates PPO and directed hypergraph module to extract the spatio-temporal attributes of the road networks. DHG-PPO enables multiple agents to ingeniously interact through the dynamical construction of hypergraph. The effectiveness of DHG-PPO is validated in terms of average travel time and throughput against state-of-the-art baselines through extensive experiments.
Index Terms:
Traffic signal control, directed hypergraph learning, multi-agent system, deep reinforcement learning.
I Introduction
Traffic Signal Control Systems (TSCs) are essential components of contemporary transportation infrastructure, playing a vital role in managing and coordinating the flow of traffic across road networks. Traditional traffic signal control methods rely on presetting light timings based on historical traffic patterns, periodically adjusted to manage vehicle flow through intersections[1]. However, the substantial rise in vehicle numbers has introduced challenges such as a lack of adaptability to dynamic traffic conditions and rudimentary delay coordination among traffic lights within the network. As a result, these systems fail to effectively mitigate traffic congestion throughout the road networks[2].
In recent years, the integration of artificial intelligence (AI) and Internet of Things (IoT) technologies has been explored to enhance intelligent traffic light control systems. Leveraging IoT sensor data from nodes across the road networks, coupled with advanced big data analytics, these intelligent systems can autonomously adjust the timing of traffic signal phases to reflect real-time traffic conditions. This approach significantly improves the responsiveness and efficiency of traffic management[3], which enhances road network throughput and alleviates the pressure of traffic congestion. Additionally, a line of research [4, 5, 6] focuses on using Deep Reinforcement Learning (DRL) to tackle the complex situation of road networks, aiming to model the traffic environment and achieve intelligent traffic light control. By using reward feedback, DRL progressively learns to make optimal decisions for signal light phase selections within the road networks. Some classic DRL algorithms such as Deep Q-Network (DQN) and Asynchronous Advantage Actor Critic (A3C) have been implemented in the traffic control use case with promising results[7, 8]. However, these methods primarily focus on optimizing individual intersections (single-agent) without considering the interactions between intersections, which may result in local optimal solutions [9]. In contrast, multi-agent reinforcement learning enables multiple traffic lights to collaborate and learn cooperative strategies. For instance, the work in [10] merged the cooperative vehicle infrastructure system into a value-based multi-agent reinforcement learning model and designed a spatial discount reward, facilitating collaborative strategy optimization among traffic lights. Nonetheless, due to the complexities of the large action space dimensions, achieving the maximum reward value is still challenging.
On the other hand, recent advancements in Graph Neural Networks (GNNs) have improved the representation embeddings of traffic signal control nodes by utilizing the inherent topological structure of the road network [11]. Inspired by the idea of incorporating the reward feedback mechanism into the learning process, some recent studies have employed spatio-temporal graph neural network methods combined with multi-agent reinforcement learning to reduce vehicle travel time and alleviate congestion in the road network, yielding promising results[12, 13, 14]. However, these approaches do not comprehensively account for the impact of spatio-temporal traffic flow between adjacent intersections. Meanwhile, compressing complex relationships into pairwise relationships loses high-order spatial-temporal information, posing challenges in handling real-world dynamic road network situations. To address this issue, hypergraph learning offers a promising approach to capture the complex spatio-temporal relationships among neighboring intersections [15]. In contrast to the conventional graph in pairwise style, hypergraphs utilize hyperedges to connect more than two nodes to form hyperedges, by which provides distinct benefits in modeling complex relationships among multiple nodes and enable comprehensive higher-order network analysis [16, 17].
To this end, we propose a novel hypergraph-based multi-agent DRL framework for mining high-order hidden correlations among traffic lights in the road network. Moreover, we introduce Proximal Policy Optimization (PPO), which employs a simple yet effective clipping mechanism that restricts the magnitude of policy updates, preventing significant policy changes during the learning process. In large-scale complex environments, PPO is an effective solution among RL algorithms [18]. The main contributions of this paper are as below: (1) Framework: We propose a hypergraph-based multi-agent reinforcement learning framework. Agents are deployed at each intersection to coordinate traffic flow conditions between multiple intersections. Interaction between agents is realized by dynamically constructing directed hypergraphs. Combined with Multi-Agent Proximal Policy Optimization (MA-PPO), signal light decision optimization between multiple intersections strikes a balance between network-level and single-intersection-level. (2) Directed hypergraph learning: The directed hypergraph module is introduced to extract the directed spatial attributes of the road network according to the driving direction of traffic flow. Spatial-temporal correlations among multiple node groups are captured through hyperedges generated by a couple of dynamical headset and tailset, which contribute to updating node embeddings via the multi-head attention mechanism. (3) Experiments: We conduct rigorous experiments based on real-world road network datasets. The obtained results validate the effectiveness of our proposal compared with several state-of-the-art methods in terms of average travel time and throughput.
II Related Work
II-A Reinforcement Learning based methods for TSC
Traditional TSC methods use deterministic offline strategies such as phase switching with predefined rules [19, 20, 21]. In contrast, reinforcement learning methods seek to adjust the signal light switching strategy adaptively based on the real-time situation at the intersection[22]. Although these DRL methods can create effective optimization strategies within individual intersections, the collaborative interaction between intersections is neglected. Therefore, when considering the overall road network, it may become trapped in local optimal situations. To mitigate this issue, max-pressure control [23, 24, 25], independent Advantage Actor-Critic (A2C) [26], and policy gradient method [27] are employed with multi-agent DRL schema, leading to a collaboration of various intersections for the global traffic optimization. However, the aforementioned studies primarily focus on extracting information from the first-order neighbors of each intersection, which overlooks the potential essential high-order information across different intersections.
II-B Integration of Reinforcement Learning and Graph Learning
Due to the inherent topological properties of road networks, our proposed framework is closely related to the recent advances of GNN-based studies[13, 12, 28, 29]. Therein, Colight, proposed by Wei et al. [13], utilizes Graph Attention Networks (GAT) combined with parameter-sharing to balance the influences among adjacent intersections. Nishi et al. [12] demonstrate a model that uses Graph Convolutional Networks (GCN) to learn representations among multiple adjacent nodes. Although these two methods utilize the message-passing strategy in graph learning to absorb the representation information of adjacent nodes, the traffic flow direction on the graph are not incorporated. Comparatively, we present a directed hypergraph-based framework to enhance high-order correlation extraction among multiple nodes in the road networks.
III Methodology
III-A Problem definition
This paper frames the control of traffic signal lights as a multi-agent reinforcement learning problem, deploying an intelligent agent at each intersection to manage the switching of signal light phases. To model the relationship between the traffic state and control action, it is reasonable to simplify the traffic signal control as the Markov decision process, and the effectiveness is supported by numerous empirical studies[5]. The problem is characterized by the following components:
State space $\mathcal{S}$ and Observation space $\mathcal{O}$: At time $t$, each agent $i$ can only access a local observation $o_{i}^{t}$ from local system state $s_{i}^{t}$. $o_{i}^{t}$ consists of agent information, its current phase represented by a one-hot vector, the number of vehicles on each lane connected with the intersection.
Action $\mathcal{A}$: We assume each intersection contains four signal phases: Phase 1 (going straight from West and East), Phase 2 (going straight from North and South), Phase 3 (turning left from West and East), Phase 4 (turning left from North and south). At time $t$, each agent $i$ selects one of four signal light phases as an action $a_{i}^{t}$.
Reward $\mathcal{R}$ and Policy $\pi$ : After agent $i$ select the action, it obtains a reward $r_{i}^{t}$ by reward function $\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}\rightarrow\mathcal{R}$. We define the immediate reward for agent $i$ as $r_{i}^{t}=-\sum_{l}u_{i,l}^{t}$, where $u_{i,l}^{t}$ is the queue length on the approaching lane $l$ at time $t$. Agent $i$ chooses an action $a$ based on a certain policy $\mathcal{O}\times\mathcal{A}\rightarrow\pi$, aiming to maximize the total reward [13].
III-B Framework
We propose a novel framework based on MA-PPO, in which PPO is an on-policy gradient method for reinforcement learning [30]. Meanwhile, to effectively capture the spatio-temporal correlations among multiple sets of intersections, we introduce directed hypergraphs to enhance coordination among multiple agents deployed at intersections. In Figure 1, the module of directed hypergraph learning is consistent with the Spatial-Temporal Hypergraph Learning module proposed in [31].
Figure 1: An overview of our proposed framework.
III-B1 Directed Hypergraph
Due to the complex spatio-temporal nature of traffic flow, intersections exhibit intricate spatio-temporal correlations. Unlike edges in standard graphs that only describe relationships between two nodes, directed hyperedges in directed hypergraphs can capture the complex spatio-temporal correlations between two sets of nodes, i.e., head node set and tail node set [16].
Inspired by [31], we adopt the dynamic construction of directed hypergraphs to capture the spatio-temporal correlations between multiple sets of intersections by
$h_{i}^{t}=\sigma(o_{i}^{t}W_{e}+b_{e})$ (1)
where $o_{i}^{t}\in\mathbb{R}^{d}$ is the observation of agent $i$ at time $t$. $W_{e}\in\mathbb{R}^{d\times l}$ and $b_{e}\in\mathbb{R}^{l}$ are trainable weight matrix and bias vector. $\sigma(\cdot)$ is the ReLU function. $h_{i}^{t}\in\mathbb{R}^{l}$ represents the traffic information of the $i$-th intersection at time $t$. ${H}_{\mathcal{G}}^{t}=\{h_{1}^{t},...,h_{N}^{t}\}\in R^{N\times l}$ is the initial node embedding of the entire road network.
Then, we capture the complex spatio-temporal correlations between nodes in the road network by dynamically constructing spatial hyperedges and temporal hyperedges. The spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$ can be generated based on the reconstruction of the master node $\dot{v}^{t}_{i}\in\emph{V}^{t}$ and the spatial candidate slave node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\emph{V}^{t},v\neq\dot{v}^{t}_{i}\right\}$, which is denoted as:
${c}_{spa}\left(\dot{v}^{t}_{i}\right)=\|\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)\cdot\theta_{spa}-\emph{p}^{spa}_{\dot{v}^{t}_{i}}\cdot\emph{H}_{\mathcal{G}}^{t}\left(\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)\right)\|_{2}~{}$ (2)
where ${c}_{spa}\left(\dot{v}^{t}_{i}\right)$ denotes the spatial reconstruction error. $\left\|\cdot\right\|_{2}$ denotes the $l2$ norm of the vector. $\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)$ and $\emph{H}_{\mathcal{G}}^{t}\left(\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)\right)$ are node feature matrices of the master node and the spatial candidate slave node set respectively. $\theta_{spa}$ is a trainable projection matrix when generating the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$. $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{\left(N-1\right)}$ is the trainable reconstruction coefficient vector. $S^{spa}\left(\dot{v}^{t}_{i}\right)$ is denoted as:
$S^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right),\emph{p}^{spa}_{\dot{v}^{t}_{i}}\left(v\right)>0\right\}$ (3)
Based on the trainable vector $\emph{p}^{spa}_{\dot{v}^{t}_{i}}$, we filter the slave nodes to construct the tail node set $\mathcal{H}_{T}$ in a more flexible manner. We group master nodes that share the same slave node set to form the head node set $\mathcal{H}_{H}$. Finally, we form directed hyperedges composed of head and tail node-sets. Similarly, temporal slave nodes are encircled from $\emph{V}^{t-1}$ to form the temporal hyperedge. Overall, the loss of hyperedges generation in one timestamp is defined as:
$\begin{split}\mathcal{L}_{recon}\!\!=\!\!\!\!\sum_{i=\left[1,...,N\right]}&\!\!\!\!\!\!\lambda\left(c_{spa}\!\left(\dot{v}^{t}_{i}\right)\!+\!c_{tem}\!\left(\dot{v}^{t}_{i}\right)\right)\!+\!\left(\|\emph{p}^{spa}_{\dot{v}^{t}_{i}}\|_{1}+\|\emph{p}^{tem}_{\dot{v}^{t}_{i}}\|_{1}\right)\\ &+{\gamma_{2}}\left(\|\emph{p}^{spa}_{\dot{v}^{t}_{i}}\|_{2}+\|\emph{p}^{tem}_{\dot{v}^{t}_{i}}\|_{2}\right)~{}\end{split}$ (4)
where $\left\|\cdot\right\|_{1}$ denotes the $l1$ norm of the vector and $c_{tem}$ denotes the reconstruction error of temporal hyperedges. $\lambda$ is the weight hyperparameter. $\gamma_{2}$ is the regularizing factor to balance $l1$ norm and $l2$ norm.
Let $\emph{M}\in R^{N\times N}$ represents the correlation between all nodes and the master node $\dot{v}^{t}_{i}$ at time $t$, with entries are as below:
$\emph{M}\left(v,e\left(\dot{v}^{t}_{i}\right)\right)=\left\{\begin{matrix}1\text{,}&v=\dot{v}^{t}_{i}\\ \emph{p}_{\dot{v}^{t}_{i}}\!\left(\!v\!\right)\!\text{,}&v\in\mathcal{H}_{T}\left(\!\dot{v}^{t}_{i}\!\right)\\ {1}/{|\mathcal{H}_{H}\left(\!\dot{v}^{t}_{i}\!\right)|}\text{,}&v\in\mathcal{H}_{H}\left(\!\dot{v}^{t}_{i}\!\right)\\ 0\text{,}&\text{otherwise}\end{matrix}\right.$ (5)
The embedding of hyperedges is aggregated by node features following:
$\emph{E}\left(e\left(\dot{v}^{t}_{i}\right)\right)=\frac{\sum_{v\in\mathcal{V}^{t}}\emph{M}\left(v,e\left(\dot{v}^{t}_{i}\right)\right)\times\emph{H}_{\mathcal{G}}^{t}\left(v\right)}{\sum_{v\in\mathcal{V}^{t}}\emph{M}\left(v,e\left(\dot{v}^{t}_{i}\right)\right)}$ (6)
We calculate multi-head attention between the master node and the two types of hyperedge ($e_{spa}\left(\dot{v}^{t}_{i}\right),e_{tem}\left(\dot{v}^{t}_{i}\right)$). The calculation process for the weights of the two types of hyperedge is similar, and thus we obtain spatial hyperedge attention by:
$att^{h}\left(\dot{v}^{t}_{i}\right)=\frac{\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)\cdot Q\text{-}Lin^{h}\cdot\Theta_{spa}^{att}\cdot\emph{K}_{spa}^{h}\left(\dot{v}^{t}_{i}\right)^{T}}{\sqrt{d}}$ (7)
where
$\emph{K}_{spa}^{h}\left(\dot{v}^{t}_{i}\right)=\emph{E}\left(e_{spa}\left(\dot{v}^{t}_{i}\right)\right)\cdot K\text{-}Lin^{h}_{spa}$ (8)
Here, for the $h$-th attention head $att^{h}\left(\dot{v}^{t}_{i}\right)$, $Q\text{-}Lin^{h}\in\mathcal{R}^{d\times\frac{d}{K}}$ is linear transformation matrix, and $K\text{-}Lin^{h}_{spa}$ is a trainable matrix with the same dimensions. $K$ is the number of attention heads. $\Theta_{spa}^{att}\in\mathcal{R}^{\frac{d}{K}\times\frac{d}{K}}$ is trainable weight matrix. The $h$-th temporal attention $att^{h}\left(\dot{v}^{t}_{i}\right)$ is calculated in a similar manner. The weight of the spatial hyperedge $w^{h}_{spa}\left(\dot{v}^{t}_{i}\right)$ and the temporal hyperedge $w^{h}_{tem}\left(\dot{v}^{t}_{i}\right)$ are calculated by $softmax$ normalization to $att^{h}\left(\dot{v}^{t}_{i}\right)$ and $att^{h}\left(\dot{v}^{t}_{i}\right)$.
$\begin{split}\emph{Q}_{\emph{V}^{t}}\left(\dot{v}^{t}_{i}\right)=\|_{k=1}^{K}\Bigl{(}w^{h}_{spa}\left(\dot{v}^{t}_{i}\right)\times\emph{K}_{spa}^{h}\left(\dot{v}^{t}_{i}\right)\\ +w^{h}_{tem}\left(\dot{v}^{t}_{i}\right)\times\emph{K}_{tem}^{h}\left(\dot{v}^{t}_{i}\right)\Bigl{)}\end{split}~{}$ (9)
[…]
III-B2 Multi-agent Deep Reinforcement Learning
PPO-Clip[30] modifies the surrogate objective by clipping the probability ratio, to penalize changes to the policy that move $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ away from 1. The policy loss function is defined as:
$L^{clip}_{\pi}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (10)
where $\epsilon$ is a hyperparameter indicating the range for clipping. $\hat{A}_{t}$ is an estimator of the advantage function at time $t$ [30].
To realize the procedure in an end-to-end fashion, the loss function used in the training process is denoted as:
$\mathcal{L}_{\textit{HG}}=\beta\mathcal{L}_{recon}+\left(1-\beta\right)\textit{MSE}\left(\textit{MLP}\left(\emph{Q}_{\mathcal{G}^{t}}\right),\emph{y}^{t}\right)~{}$ (11)
where MSE is the mean squared error function. $\beta$ is a weight hyperparameter to balance the impact of reconstruction loss and Q-value loss. $y^{t}$ is the target Q-value.
IV Experiments
We establish a simulation testbed based on CityFlow [32], which is an open-source traffic simulator that supports large-scale traffic signal control. Two real-world datasets $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$ [13] are adopted in the experiment.
We introduce the following five methods as baseline: Fixed-time [19]: Assigning a constant cycle length with a predetermined green ratio distribution across all phases. MaxPressure [21]: Controlling traffic signals by alleviating congestion on the upstream and downstream queue length with the max pressure. GCN [12]: Using graph convolution neural network to extract traffic features among multiple intersections. PressLight [24]: Integrating pressure into the state and reward design of the reinforcement learning model to solve the multi-intersection signal control problems. MPLight [25]: A decentralized network level traffic signal control RL algorithm with parameter sharing which enables large scale application based on Presslight. CoLight [13]: Using a graph attention network to enhance coordination and decision-making of multi-intersection traffic signal control.
TABLE I: Performance of average travel time and throughput
  Method Average travel time(s) Throughput
$D_{\textit{Hangzhou}}$ $D_{\textit{Jinan}}$ $D_{\textit{Hangzhou}}$ $D_{\textit{Jinan}}$
  Fixed-time 718.29 814.11 1989 3475
MaxPressure 407.17 343.90 2664 5613
  GCN 709.70 596.67 2760 6024
PressLight 416.73 311.39 2832 6116
MPLight 319.71 304.93 2927 6125
CoLight 302.37 309.93 2930 6116
DHG-PPO 292.56 277.94 2930 6158
Table I summarizes the performance of our proposed DHG-PPO against baselines in real-world datasets. The results demonstrate that DHG-PPO outperforms the baselines in minimizing average travel time (ATT) and maximizing throughput metrics. Here, ATT is defined as the average time spent by all vehicles between entering and leaving the area, while throughput is defined as the number of trips completed by vehicles throughout the simulation.
For $D_{\textit{Jinan}}$, DHG-PPO achieves a maximum reduction in ATT of 8.85% compared to the second optimal solution MPLight. Additionally, the maximum enhancement of throughput over the second-best MPLight reaches 0.53%. Traditional control methods (Fixed-time and MaxPressure) rely on predetermined strategies to regulate traffic signals, which results in unsatisfactory performance in terms of ATT and throughput. In contrast, GCN and CoLight utilize GNNs to enable interactions between multiple intersections to improve overall efficiency. However, in real-world datasets, the dynamic changes in traffic flow create higher-order correlations between intersections that traditional graphs cannot fully capture. PressLight and MPLight incorporate the concept of Maxpressure at each intersection but do not account for the dynamic changes in traffic flow that introduce complex directed spatio-temporal attributes to the entire road network.
(a) $D_{\textit{Hangzhou}}$
(b) $D_{\textit{Jinan}}$
Figure 2: Ablation results on two real-world datasets.
Figure 2 plot the results of albation study. Here, MA-PPO utilizes a different graph module (GAT) from our proposed model to capture correlations between intersections. In contrast, MA-SAC and DHG-SAC incorporate the SAC algorithm to validate the effectiveness of hypergraphs. The results show that DHG-PPO and DHG-SAC converge more rapidly and smoothly, which validates our motivation that hypergraphs offer superior performance in capturing correlations between multiple sets of intersections.
Figure 3 investigates the impact of the hyperparameter $\beta$ in Equation 11. The results indicate that the optimal values of $\beta$ are 0.01 and 0.3 for the $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$, respectively. As the value of $\beta$ increases, the proportion of the loss associated with dynamically constructing directed hypergraphs rises, which deteriorates the performance of ATT.
(a) $D_{\textit{Hangzhou}}$
(b) $D_{\textit{Jinan}}$
Figure 3: Impact of hypergraph loss coefficient $\beta$.
V Conclusion
In this paper, we propose a multi-agent reinforcement learning model DHG-PPO, which incorporates a directed hypergraph module to capture the correlation in road traffic by dynamically constructing spatio-temporal hyperedges. Extensive experimental results demonstrate the superiority of DHG-PPO against the baselines in traffic signal control. Our future work is to enhance the robustness of our method by studying the data cleaning schemes to mitigate the impact of missing or noisy existing in the collected data.